OpenStudy (anonymous):

5. which of ( 3,5 ) , (4,6 ) ,( 5,7 ) , and ( 6,8 ) are solutions to y = x + 2 ? ( 1point ) A. ( 3,5 ) and ( 4,6 ) b. ( 4,6 ) and ( 5,7 ) c. none d. all

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

y = x + 2 Plug in (3, 5): 5 = 3 + 2 Add: 5 = 5 Correct. Plug in (4, 6): 6 = 4 + 2 6 = 6 Correct. Plug in (5, 7): 7 = 5 + 2 7 = 7 Correct. Plug in 6, 8: 8 = 6 + 2 8 = 8 Correct. So the answer is all of them.